40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 800     Entry time: Tue Aug 5 17:56:23 2008
Author: Alberto 
Type: Configuration 
Category: General 
Subject: SRM and PRM inspection 
Yoichi, Koji, Rana, Steve, Alberto

Today we opened the BSC to inspect the optics, and in particular the SRM and PRM.
We found that one of the side magnets of the SRM was broken and a piece of it fell and got stuck to the LR magnet.
We removed the LR OSEM and took off the broken part with tweezers. Since we couldn’t replace the magnet on the side,
we decided to just switch the OSEM to the other side were a second magnet was available. Then we centered the OSEMs.
Using the optical levers we aligned both the ITMX and the SRM so that now we have to center again the OSEMs on both.

The PRM was visibly tilted and it was out of the range of the OSEMs. To try to fix the tilt we lift it up a little
with the screws on the bottom and pushed it with the third screw on top. That had the effect of making the mirror
tilt to the opposite direction. We looked at the wires (see attached picture) and it seemed centered on the side
of the mirror.

Tomorrow we are going to reset the OSEMs on ITMX and SRM and then we’re going to try to fix the tilt on PRM.
Attachment 1: IMG_1434.JPG  2.432 MB  Uploaded Wed Aug 6 11:57:13 2008  | Hide | Hide all
IMG_1434.JPG
Attachment 2: IMG_1456.JPG  2.461 MB  Uploaded Wed Aug 6 11:57:32 2008  | Hide | Hide all
IMG_1456.JPG
ELOG V3.1.3-